# TabSTAR: A Foundation Tabular Model With Semantically Target-Aware Representations

### v1

@@ F.2 Head-to-head comparisons
CARTE CatB CatB-T RF TabPFN XGB XGB-T
Classification 93.3±5.2 73.6±7.3 67.9±7.8 88.6±5.3 57.3±9.3 89.3±5.1 71.4±7.5
Regression 35.7±5.9 33.6±4.9 32.2±4.8 66.9±4.9 40.9±5.2 69.4±4.8 30.3±4.8
Table 22: Win rates of TabSTAR-U (above 10K) against baselines. Win rate with 95% CI.
CARTE CatB-T CatB-T-U TabPFN TabSTAR XGB-T XGB-T-U
Classification 98.6±2.8 86.0±6.8 67.0±9.3 75.7±10.1 94.0±4.7 85.0±7.0 71.0±8.9
Regression 63.5±7.5 53.0±6.9 22.0±5.8 61.1±7.0 82.0±5.3 59.5±6.8 28.5±6.3
F.3 Running Times and Compute Information
In this section we present the running times of the different models (see Table 23). We focus on the condition with up to 10,000 examples, excluding CatBoost-Tuned and XGBoost-Tuned as they run with a 4 hours budget using 8 CPUs from AMD EPYC 7742 64-Core Processor .
Compute Information
TabSTAR runs on NVIDIA RTX A40 with 48GB memory. TabPFN-v2 runs directly on their API client. CatBoost, XGboost and Random Forest run using Intel(R) Core(TM) i9-10920X CPU @ 3.50GHz CPU cores. CARTE use inferior GPUs of type NVIDIA GeForce GTX 1080 Ti with 11GB memory, to allow for higher parallelization of their incredibly slow runs.
@@ F.3 Running Times and Compute Information
TabSTAR’s average running time for a downstream task ranges from 30 seconds (C04) to 7,692 seconds (R24). We observe that these running times are significantly lower than 14,400 seconds (4 hours), used for CatBoost-Tuned and XGBoost-Tuned, although a single run of any tree model without 5-fold cross-validation is faster. The running times of TabPFN-v2 are considerably better. However, since they use ICL, they are limited to run only up to 10,000 examples and their inference time is just as costly as their fitting time. For CARTE, while their runs use inferior hardware and limited parallelization, the reported times are for a single run. In practice, they run 6 times per dataset split because of their lack of default hyperparameters, making it a poor choice for scale.
@@
ID CARTE CatBoost RF TabPFN-v2 TabSTAR XGBoost
C01 13,341 44 19 42 302 31
C02 420 346 1,145 491
C03 7,297 38 19 61 324 27
C04 11 6 9 30 9
C05 1,227 10 8 14 68 9
C06 3,068 68 31 112 1,126 63
C07 6,559 61 34 83 685 52
C08 4,631 40 41 49 681 49
C09 20 21 134 544 31
C10 22 22 30 519 23
C11 2,428 19 12 33 252 18
C12 4,411 169 16 765 39
C13 1,774 31 27 58 929 37
C14 500 140 5,254 308
R01 9,439 139 446 1,987 94
R02 9 16 54 188 6
R03 2,174 30 103 357 553 24
R04 10 10 84 115 7
R05 10,639 28 26 62 158 15
R06 3,185 22 43 70 307 13
R07 4,637 22 109 112 352 28
R08 4,252 57 366 396 1,136 58
R09 8,239 7 22 55 198 5
R10 1,505 28 37 33 168 21
R11 11 18 36 99 8
R12 13 22 78 199 9
R13 2,299 24 54 38 639 21
R14 6,396 28 30 167 772 15
R15 25 51 66 223 22
R16 13,380 34 118 156 798 35
R17 3,247 94 478 204 270 143
R18 9,663 34 631 309 341 32
R19 11 14 66 101 6
R20 6,173 22 144 154 637 20
R21 2,215 14 33 25 177 14
R22 16 21 17 149 15
R23 14,401 15 51 58 167 10
R24 25,903 90 170 314 7,692 106
R25 2,404 50 943 261 786 54
R26 636 10 11 13 71 8
R27 2,615 31 59 40 304 22
R28 12,265 45 163 128 389 46
R29 11 10 32 138 5
R30 2,477 9 18 24 45 7
R31 5,102 41 187 240 451 39
R32 8 11 13 32 7
R33 4,512 28 100 136 539 30
R34 26 24 18 61 12
R35 4,816 30 55 48 396 28
R36 15,389 18 81 79 949 23
@@ G.1 Evaluation Datasets for Analysis
All the experiments are evaluated over 20 datasets from the benchmark in Appendix D. Each experiment reports the performance with AUROC for classification and $R^{2}$ for regression. For conciseness, each tables reports both regression and classification tasks, distinguishable by their ID.
@@ G.2 The Role of the Encoder Unfreezing (Q1)
Table 24 shows the results for each variant of the experiment presented in §6. It is evident that unfreezing the textual encoder yields a significant performance across datasets. Furthermore, while finetuning a single layer gives a significant boost, it underperforms compared to 6 unfrozen layers.
G.3 The Effect of Pretraining (Q2)
While LoRA [41] is a very efficient technique, its performance was much worse for a randomly initialized model. Therefore, we perform full finetuning of the non-pretrained model, as explained in Appendix B.2. Table 25 shows the dataset level results. It is evident that for most of the datasets, improvement is observed when scaling, with the 256 datasets variant winning for almost all datasets.
G.4 Numerical Verbalization (Q3)
We show the full results for the experiment in §6, with Table 26 illustrating the verbalizations in each variant. Note that we do not include an exact value verbalization, since it would increase the number of unique text inputs and place extra memory demands. The two variants which integrate numerical information into the verbalization dominate the experiment, although the improvement seems to be marginal for some datasets. Interestingly, some datasets significantly underperform, with the R27 dataset completely failing the task. The addition of the quantile information on top of the bin seems to have limited impact, although marginally winning on the average performance.

### v2

@@ F.3 Head-to-head comparisons
Model Classification Regression
CARTE 93.3 ± 5.2 35.7 ± 5.9
CatBoost 80.7 ± 6.6 33.6 ± 4.9
CatBoost-Tuned 75.0 ± 7.2 32.2 ± 4.8
LightGBM 92.1 ± 4.5 49.7 ± 5.2
LightGBM-Tuned 90.7 ± 4.8 45.3 ± 5.1
RandomForest 95.7 ± 3.4 66.9 ± 4.9
RealMLP 82.1 ± 6.4 51.4 ± 5.2
RealMLP-Tuned 83.6 ± 6.2 49.4 ± 5.2
TabDPT 73.6 ± 7.3 66.7 ± 4.9
TabICL 70.0 ± 7.6
TabM 64.3 ± 8.0 39.2 ± 5.0
TabM-Tuned 63.3 ± 8.0 37.4 ± 5.0
TabPFN-v2 66.4 ± 8.9 40.9 ± 5.2
XGBoost 96.4 ± 3.1 69.4 ± 4.8
XGBoost-Tuned 78.6 ± 6.8 30.3 ± 4.8
Table 24: Win rates of TabSTAR-Unlimited (above 10K) against baselines. Win rate with 95% CI.
Model Classification Regression
CARTE 98.6 ± 2.8 63.5 ± 7.5
CatBoost-Tuned 96.0 ± 3.9 53.0 ± 6.9
CatBoost-Tuned-Unlimited 77.0 ± 8.3 22.0 ± 5.8
LightGBM-Tuned 98.0 ± 2.8 69.5 ± 6.4
LightGBM-Tuned-Unlimited 92.0 ± 5.3 44.0 ± 6.9
RealMLP-Tuned 97.0 ± 3.4 72.0 ± 6.2
RealMLP-Tuned-Unlimited 86.0 ± 6.8 38.9 ± 7.0
TabDPT 85.0 ± 7.0
TabICL 85.0 ± 7.0
TabICL-Unlimited 82.0 ± 7.6
TabM-Tuned 76.8 ± 8.4 65.2 ± 6.7
TabM-Tuned-Unlimited 53.0 ± 9.8 22.5 ± 5.8
TabPFN-v2 90.0 ± 7.1 61.1 ± 7.0
TabSTAR 94.0 ± 4.7 82.0 ± 5.3
XGBoost-Tuned 95.0 ± 4.3 59.5 ± 6.8
XGBoost-Tuned-Unlimited 81.0 ± 7.7 28.5 ± 6.3
F.4 Running Times and Compute Information
Hardware
All baselines are evaluated using a single NVIDIA A100-SXM4 GPU with 40GB memory, and 8 CPU cores of type AMD EPYC 7742 64-Core Processor. The only exclusions are TabPFN-v2, which runs on its API version, and CARTE, for which we employ several NVIDIA GeForce GTX 1080 Ti GPUs with 11GB memory, as their high latency and lack of default hyperparameters forced us to rely on more accessible hardware.
@@ F.4 Running Times and Compute Information
Table 25 presents the average running times of the different models per dataset, with up to 10,000 examples, excluding the tuned variants, TabPFN-v2 and CARTE. TabSTAR’s average running time for a downstream task ranges from 59s (C04) to 9,332s (R24). Although higher than a single run of other baselines, these times remain far below the 14,400s (4 hours) typically required for their tuned counterparts. TabICL achieves considerably faster running times; however, their ICL-based approach limits them to smaller datasets, and inference is as costly as training (see §5). TabPFN-v2, though unreported, is subject to the same limitation.
@@
ID CatBoost LightGBM RandomForest RealMLP TabICL TabSTAR XGBoost
C01 116 35 48 148 37 397 36
C02 251 57 75 179 58 1,135 58
C03 106 51 61 147 49 553 48
C04 29 17 19 25 20 59 18
C05 26 13 16 63 13 132 13
C06 170 79 83 126 77 938 81
C07 163 78 98 186 80 1,044 78
C08 89 72 86 180 74 1,057 69
C09 33 25 33 128 29 414 25
C10 49 35 43 126 36 572 35
C11 43 22 29 121 20 521 20
C12 614 39 37 145 28 729 36
C13 70 55 68 163 55 551 50
C14 1,938 307 333 418 311 4,068 329
R01 488 161 1,204 368 2,908 161
R02 28 10 45 128 283 10
R03 111 62 282 171 592 61
R04 21 9 17 138 240 7
R05 65 20 68 90 366 23
R06 60 24 119 129 501 23
R07 66 39 285 145 673 44
R08 213 105 951 222 1,448 112
R09 12 8 56 115 604 7
R10 102 45 104 72 255 47
R11 26 10 45 98 197 10
R12 34 10 42 133 335 11
R13 78 44 145 94 484 46
R14 67 24 91 159 1,296 27
R15 38 18 102 132 390 17
R16 108 76 316 193 898 80
R17 62 54 1,110 160 261 53
R18 124 74 1,705 201 553 68
R19 22 13 38 107 198 12
R20 70 52 401 176 705 49
R21 53 29 86 62 222 30
R22 57 31 62 52 171 33
R23 41 19 137 140 281 18
R24 461 396 650 520 9,332 392
R25 191 104 2,583 223 750 110
R26 29 12 22 30 205 11
R27 103 49 164 92 334 49
R28 175 92 460 158 647 97
R29 34 8 21 100 308 7
R30 26 14 47 56 114 14
R31 148 78 505 200 797 77
R32 20 11 23 31 82 11
R33 89 55 271 161 608 47
R34 66 22 66 50 122 24
R35 99 50 146 101 411 50
R36 58 38 223 146 823 39
F.5 TabSTAR’s Cost Analysis by Number of Features
While §5 reports average computational costs, these values vary considerably across datasets, depending on factors such as the number of examples, the feature count, and the amount of textual fields. In this subsection, we examine how TabSTAR’s computational cost scales with the number of features.
Table 26 summarizes both latency and memory consumption of TabSTAR, aggregated over a group of up to 10 features, a group of 30-50 features and a group of more than 100 features. These datasets were randomly sampled from the pretraining corpus; consequently, the reported times may differ from those observed on the benchmark datasets. We observe that latency can increase by a factor of 5, and GPU memory consumption by a factor of 4. This behavior stems from the transformer’s architecture, whose computational complexity is quadratic in the sequence length. This raises concerns for training TabSTAR over datasets with hundreds of features, and calls for improvements to expand to datasets of thousands of features.
Additionally, TabSTAR’s memory consumption in classification tasks is affected by the number of classes due to its target-aware tokens. We assume that for TabSTAR, the effective number of features equals the actual number of features plus the number of classes. For highly multiclass datasets, this can become a limiting factor.
Table 26: Cost Analysis for TabSTAR per number of features. Median training and inference times and peak memory usage on GPU and CPU, aggregated across 45 datasets.
Train Inference
Feature Group Time (s) GPU (GB) CPU (GB) Time (s) GPU (GB) CPU (GB)
Up to 10 85.6 1.9 4.4 0.5 1.6 4.4
30-50 228.8 3.1 4.5 1.0 1.7 4.4
100+ 417.9 7.8 4.5 2.5 2.3 4.4
F.6 Number of Trials for Tuned models
The tuned models are optimized with a budget of 4 hours using 8 CPU cores. Each trial is executed on a single core, ensuring that at least one trial could be completed within the allocated time. Table 27 presents the number of hyperparameter trials per dataset.
Table 27: Average number of hyperparameter trials per dataset. Tuned models: CatBoost (CTB), LightGBM (LGB), RealMLP (MLP), XGBoost (XGB). Unlimited models are marked with a ’!’.
Dataset CTB CTB! LGB LGB! MLP MLP! XGB XGB!
C01 27.1 24.4 770.2 812.0 56.0 26.7 97.7 90.0
C02 18.1 9.4 417.5 51.3 31.9 8.0 47.2 19.6
C03 18.0 15.8 312.0 338.1 37.4 25.2 42.4 39.5
C04 105.8 11283.8 642.1 458.5
C05 122.1 3373.3 98.3 228.3
C06 18.1 406.9 67.7 44.6
C07 17.3 18.5 335.6 376.6 31.8 23.7 53.4 37.7
C08 101.1 67.8 2911.4 2394.6 43.1 37.3 508.3 566.1
C09 169.0 81.3 3969.5 475.7 50.5 9.2 883.9 207.2
C10 73.7 1598.1 55.4 165.2
C11 58.3 48.5 812.4 863.4 44.8 26.0 124.1 96.3
C12 9.9 8.5 207.2 34.4 35.3 8.2 41.8 17.3
C13 85.8 27.4 1848.3 329.6 40.8 8.7 234.6 105.7
C14 8.6 8.4 50.9 36.3 14.9 9.7 13.7 11.6
R01 14.2 12.2 278.5 169.6 11.3 32.8 15.4
R02 138.9 118.6 4463.0 5214.6 48.0 30.7 742.4 1054.5
R03 55.5 64.2 1033.6 1265.9 35.8 30.0 94.3 97.8
R04 123.8 31.6 4309.9 714.9 44.1 8.7 1324.0 364.8
R05 78.8 4428.5 87.9 307.0
R06 88.8 2305.3 47.5 232.2
R07 109.9 58.8 1370.0 344.1 42.0 8.5 145.6 72.5
R08 30.5 30.1 568.6 122.9 21.4 10.4 65.7 47.1
R09 194.6 208.7 6391.0 5975.1 50.4 31.2 589.0 680.6
R10 53.7 2806.9 135.9 241.8
R11 142.9 4669.1 58.7 662.6
R12 126.8 45.6 2990.9 814.7 46.9 8.6 385.8 178.0
R13 80.9 2814.6 91.2 204.8
R14 59.9 67.6 1725.3 1542.6 39.5 15.2 200.6 177.0
R15 107.4 69.2 2896.1 1177.4 52.5 14.0 337.5 242.7
R16 65.8 46.3 1446.3 663.7 34.9 10.9 160.9 104.9
R17 79.7 44.6 1517.3 316.3 45.3 9.1 254.0 104.4
R18 39.1 43.8 1572.4 1041.4 29.4 16.6 258.8 144.5
R19 117.5 4046.2 63.1 1063.2
R20 62.6 89.1 1611.5 2176.4 36.4 30.1 250.5 258.8
R21 101.6 3284.9 122.3 159.4
R22 67.8 3584.3 259.9 127.2
R23 91.8 105.4 3179.2 2773.0 53.4 27.6 420.1 314.5
R24 52.4 38.1 867.2 417.7 35.2 10.4 104.7 60.3
R25 40.4 33.5 806.9 357.6 23.6 10.8 122.4 72.0
R26 166.5 12567.7 279.4 625.0
R27 53.9 2034.2 85.3 155.5
R28 40.9 670.8 37.5 59.3
R29 135.3 7615.2 63.8 1024.3
R30 194.3 5275.3 114.8 431.7
R31 36.6 39.5 774.5 504.3 26.4 12.6 90.3 54.1
R32 143.2 6605.7 191.7 337.3
R33 69.3 30.8 1053.9 212.0 36.7 8.3 119.8 51.1
R34 69.1 3962.2 148.9 314.9
R35 65.8 3351.6 76.3 139.9
R36 108.0 120.2 2633.9 4634.0 49.9 48.0 289.9 228.8
@@ G.1 Evaluation Datasets for Analysis
All experiments are conducted on 20 datasets from the benchmark described in Appendix D. Each experiment reports performance using AUROC for classification and $R^{2}$ for regression. Each tables reports both regression and classification tasks, distinguishable by their ID. Furthermore, each experiment compares its variants to other models, excluding TabPFN-v2 and CARTE, which could not be executed on all datasets. Although this comparison covers only a subset of the benchmark and should not be interpreted as conclusive, the results nonetheless offer valuable reference points.
@@ G.2 The Role of the Encoder Unfreezing (Q1)
Table 28 shows the results for each variant of the experiment presented in §6. It is evident that unfreezing the textual encoder yields a significant performance boost across datasets, as seen in Figure 6: The frozen TabSTAR variant is much worse than any other baseline! Furthermore, while finetuning a single layer gives a significant boost, it underperforms compared to 6 unfrozen layers.
Figure 6: Comparison of normalized scores with 95% CIs for the encoder unfreezing experiment, comparing TabSTAR variants against baseline models in classification tasks.
@@
G.3 The Effect of Pretraining (Q2)
While LoRA [42] is a powerful technique, it can’t be applied to a randomly initialized model. Therefore, we perform full finetuning of the non-pretrained model, as explained in Appendix B.2.
Table 29 show the normalized results for classification and regression, and Table 30 shows the dataset level results. It is evident that for most of the datasets, improvement is observed when scaling, with the 256 datasets variant winning for almost all datasets. Figure 7 highlights this effect, showing that adding more datasets enables TabSTAR to outperform the baselines.
Pretraining Datasets 0 16 64 256
Classification 0.352 ± 0.086 0.450 ± 0.084 0.558 ± 0.086 0.786 ± 0.076
Regression 0.338 ± 0.073 0.395 ± 0.068 0.642 ± 0.066 0.811 ± 0.055
Figure 7: Comparison of normalized scores with 95% CIs for the pretraining effect experiment, comparing TabSTAR variants against baseline models in classification tasks.
@@ G.3 The Effect of Pretraining (Q2)
G.4 Numerical Verbalization (Q3)
We show the full results for the experiment in §6, with Table 31 illustrating the verbalizations in each variant. Note that we do not include an exact value verbalization, since it would increase the number of unique text inputs and place extra memory demands. The two variants which integrate numerical information into the verbalization dominate the experiment, although the improvement seems to be marginal for some datasets. Interestingly, some datasets significantly underperform, with the R27 dataset completely failing the task. The addition of the quantile information on top of the bin seems to have limited impact, although marginally winning on the average performance. Figure 8 presents a comparison with baseline models, highlighting the importance of incorporating richer numerical verbalizations beyond relying solely on the column name.
Figure 8: Comparison of normalized scores with 95% CIs for the numerical verbalization experiment, comparing TabSTAR variants against baseline models in classification tasks.